Find the Expected value in dollars.

There are two independent outcomes both with one-in-three chance of happening.

If one of the two happens you will make 2 dollars.

If neither happens you will loose 2 dollars.

If both happen you will make 8 dollars.

What is the expected value?
Answer #1

Probability that one of the two events happens = (1/3*2/3) + (2/3*1/3) = 4/9

Probability that neither events happens = 2/3*2/3 = 4/9

Probability that both events happens = 1/3*1/3 = 1/9

The Expected value = 4/9*2 + 4/9*(-2) + 1/9*(8) = 8/9 = 0.889 dollars